OpenStudy (anonymous):

Solve the quadratic equation. 4x2 - 30 = 0 A. ± √15 B. ± 15/√2 C. ± √15 /2 D. ± √30 /2

OpenStudy (anonymous):

x^2 = 30/4 = 15/2 x = sqrt(30)/2 should be in mathematics